People with type O-negative blood are universal donors whose blood can be safely given to anyone. Only 7.2% of the population have O-negative blood. A new, larger random sample was obtained involving 300 donors from the general population. What is the probability that more than 10% of the donors have O-negative blood?
Answer #1

Using central limit theorem,

P( < p ) = P(Z < - p / sqrt [ p( 1 - p) / n ]

So,

P( > 0.10) = P(Z > 0.10 - 0.072 / sqrt [ 0.072 ( 1 - 0.072) / 300 ] )

= P(Z > 1.88)

= 0.0301